The coefficient of static friction between the palm of Abby's hand and the softball bat is 0.58. How hard should she squeeze the bat if she wants to exert a total force of 345 N along the longitudinal axis of the bat?
Answer #1

friction force is 345 newtons

relation between friction and normal force is given by

f=mu N

= f/H = 345/0.58 = 594 .827neutons